SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (30th Jan, 2007)

Attendees: Debbie Bard, Brandon Drummond, Al Eisner, Hojeong Kim, Rahmat Rahmat, Su Dong, Mark Tibbetts.

Operations

  • Colliding beams over the weekend - Run 6 can be said to have started in earnest. Trigger data looks good so far despite high backgrounds.
  • Paged three times this week:
    • Sat day shift - one DHP node down, problem fixed itself as per DQM long-term instructions. A note will be added to JAS page to emphasise this is a known problem and does not neccessitate a page.
    • Mon owl - High backgrounds caused high deadtime. Note in short-term instructions added to remind shifter to call MCC if backgrounds are high. Also, L1 Trigger plots were empty in epics - instreucted shifter to page DAQ.
    • Mon day - Crate damage appeared, but this was swiftly followed by a TRG synch request, so the pilot ended the run and started another to fix problem.
    Note: babartrg .forward wasn't working (permissions problem?) so no babartrg or pager emails have been seen for a while. Fixed now.
  • Bad DCC and FPCM boards have been sent for repairs.
  • In the ROD tomorrow I will swap the spare BLT and GLT boards in.
  • There are some old channels in epics from the DCT era, Matt Weaver may help clean them up. The unused channels have already been removed from the archiver (thanks to Andy Salnikov!).

EMT

  • Noisy channels in EMC over the weekend showed up as hot towers in the EMT JAS plots, but the shifters were slow to page EMC experts. Note to be added to JAS page to emphasis this is a problem.
  • Mark is on the ir2 experts list contact page.
  • Jong Li mentioned there are effects in EMC plots possibly due to systematic effects of EMT. Mark will talk to him.
  • Erratum of last week's minutes - the M threshold plots do not show L3 calibrated energy but offline calibrated energy.
  • Which version of L1TOprMon to use? Recipe here may be out of date.
  • L1TOprMon Emt module crashed moose. It seems there's no check for a null pointer made wich causes a segmentation violation - module uses digis that have no equivalent in the mini and this incompatibility crashes moose. Should be fixed! Hypernews thread and following.

L3

  • L3 data for the recent runs look good.
  • There's a subtle bug in the trickle stream servers - details here.

Data Quality and Monitoring

  • Data looks good so far.
  • It seems to be impossible to make a stripchart encompassing all of Run 5. Stripcharts split into 500 runs apiece are here.

Config

  • MC truth in L1TNtupleMooseApp still sin't working. Eric Torrence has kindly said he'll look into it. In the meantime, a cut on d0 and z0 will remove some of the loopers and hopefully clear up the plots.

AOB

  • We ought to have a backgrounds expert, especially as we approach high lumi period. They should attend the Friday 8.30 MDI meetings (announced in Backgroudn hypernews) - ops manager attends at present.

This page is created by Debbie Bard (djbard@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.